Calculate Delta H for the reaction 4 NH3 (g) —-> 4 NO (g) —-> + H2O (g)
1
Expert's answer
2020-10-28T08:42:45-0400

4NH3 (g) + 5O2 (g) → 4NO (g) + 6H2O (g)

N2 (g) + O2 (g) → 2NO (g) delta H = -180.5 kJ (1)

N2 (g) + 3H2 (g) → 2NH3 (g) delta H = -91.8 kJ (2)

2H2 + O2 → 2H2O (g) delta H = -483.6 kJ (3)

Multiplying equation (1) by 2 to obtain

2N2 (g) + 2O2 (g) → 4NO (g) delta H "= 2 \times (-180.5\;kJ) = -361.0 \;kJ"

Multiplying equation (2) by 2 and then reversing it to obtain

4NH3 (g) → 2N2 (g) + 6H2 (g) delta H "= 2 \times 91.8\;kJ = 183.6 \;kJ"

Multiplying equation (3) by 3 to obtain

6H2 + 3O2 → 6H2O (g) delta H "= 2 \times (-483.6 \;kJ) = -967.2 \;kJ"

Adding the last three equations to obtain

4NH3 (g) + 5O2 (g) → 4NO (g) + 6H2O (g) delta H = -361.0 kJ + 183.6 kJ – 967.2 kJ = -1144.6 kJ

Answer: -1144.6 kJ